40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 10723     Entry time: Mon Nov 17 20:40:29 2014
Author: rana, diego 
Type: Update 
Category: IOO 
Subject: Investigating the IMC WFS situation 

We've known for years that the IMC WFS sensing chain is pointlessly bad, but until recently, we haven't thought it was worth it to fix.

There are problems in all parts of the chain:

  1. The WFS Photodetectors oscillate ~200 MHz when turned up to full gain. Diego and I confirmed this today by measuring the RF spectrum of the signals going into the WFS demod boards and seeing the oscillation change (not much) with RF gain. I recommend we switch the heads into the full gain mode (turn all of the attenuators OFF). At the moment we are operating with the 2dB and 8dB attenuators ON.
  2. The demod board has some bad gain allocation and noisy opamps.
  3. The whitening board has too much up/down of gain with noise injection along the way. And the range cannot fill up the ADC.
ELOG V3.1.3-